OpenStudy (anonymous):

Determine the dimensions for enclosing the maximum area of a rectangle if: a. The perimeter is held constant at 280 meters. Assume that the length is greater than or equal to the width. The length is_______ m. The width is ________m. b. The perimeter is held constant at P meters. The length is_______ m. The width is_________ m.

OpenStudy (anonymous):

Let L be the length and W be the width. The area A is:\[A=L·W\]and the perimeter P is:\[P=2(W+L)\rightarrow L=\frac{ P }{ 2 }-W\]Then we have:\[A=\left( \frac{ P }{ 2 } -W\right)W=\frac{ P }{ 2 }W-W^2\]In order to enclose the maximum area we need to solve:\[\frac{ dA }{ dW }=\frac{ P }{ 2 }-2W=0\rightarrow W=\frac{ P }{ 4}\rightarrow L=\frac{ P }{ 2}-\frac{ P}{ 4 }=\frac{ P}{4 }\] As\[L=W=\frac{ P }{ 4 }\], it is a square

OpenStudy (anonymous):

for P=280 we get L=W=280/4=70

OpenStudy (anonymous):

Thank you so much you were right:)

OpenStudy (anonymous):

welcome